OpenStudy (anonymous):

Find the limit. \[\lim_{x \rightarrow \infty}(\sqrt{x^2+x}-\sqrt{x^2-x})\]

myininaya (myininaya):

multiply top and bottom by the conjugate of the top

myininaya (myininaya):

that is what you have is over 1 already

myininaya (myininaya):

\[\frac{\sqrt{x^2+x}-\sqrt{x^2-x}}{1} \cdot \frac{\sqrt{x^2+x}+\sqrt{x^2-x}}{\sqrt{x^2+x}+\sqrt{x^2-x}}\]

myininaya (myininaya):

once you do that I bet you can decide what to divide top and bottom by if not let me know

OpenStudy (anonymous):

I got 1 as the numerator, is that correct?

myininaya (myininaya):

when you multiplied?

OpenStudy (anonymous):

numerator: x^2+x-x^2-x

myininaya (myininaya):

oh you forgot to distribute

myininaya (myininaya):

you have (x^2+x)-(x^2-x) right?

myininaya (myininaya):

x^2+x-x^2+x

OpenStudy (anonymous):

oh oops!

myininaya (myininaya):

by if you had what you said that would actually be 0 not 1

myininaya (myininaya):

but anyways it is neither 0 or 1 in our case it is actually just 2x on top

OpenStudy (anonymous):

Yes, I got 2x

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{2x}{\sqrt{x^2+x}+\sqrt{x^2-x}}\]

myininaya (myininaya):

do you know what you are going to choose to divide top and bottom by?

OpenStudy (anonymous):

no...

myininaya (myininaya):

ok notice if the radical wasn't there there the highest exponent on bottom is 2 so you would have said x^2 but we have a radical over the x^2 so you divide top and bottom by sqrt(x^2)

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{\frac{2x}{\sqrt{x^2}}}{\sqrt{\frac{x^2}{x^2}+\frac{x}{x^2}}+\sqrt{\frac{x^2}{x^2}-\frac{x}{x^2}}}\]

myininaya (myininaya):

now since sqrt(x^2)=x if x>0 then you can replace the top sqrt(x^2) with x if you had x was going to negative infinity you would replace sqrt(x^2) with -x instead

OpenStudy (anonymous):

Okay... I got this \[\frac{ 2 }{ 1+1 }= 2/2=1\]

myininaya (myininaya):

great work

OpenStudy (anonymous):

Thank you :)

myininaya (myininaya):

np :)